Status Report

NOAA GOES SATOPS Morning Report 1 Dec 2000

By SpaceRef Editor
December 1, 2000
Filed under ,

SATOPS MORNING REPORT – December 1, 2000

POES

No change in the status of the POES spacecraft.

NOAA-16 SBUV OV re-testing, which began at 0132Z on Thursday, was completed
at 1132Z today, December 1.

The NOAA-16 VTX2 transmitter was cycled yesterday in an attempt to restore
functionality.  The transmitter was first turned on over FCDAS during
rev 991 at 1644Z.  The transmitter was then power cycled in a warm
condition during rev 992 / W at 1809Z and then the RF switch was cycled
between VTX2 and VTX1.  During rev 993 / W at 1952Z, one additional
attempt to cycle VTX2 was made unsuccessfully subsequent to exercising
the RF switch.  Both transmitters appeared to put out an equivalently
low power signal, indicating a potential problem in the common portion
of the RF path (RF switch to the antenna).  Engineers are currently
analyzing the test data and planning the next step(s).

The NOAA-15 AVHRR was re-synchronized with the on-board information
processor (MIRP) with rephasing disabled yesterday during rev 13254 / W
at 1410Z.  Although it was anticipated that this configuration would
not effect the operational (mostly AMSU) data, engineers returned the MIRP
to internal synch during rev 13255 / F at 1547Z because of observed data
dropouts.  Further discussions with NASA and the instrument vendor,
ITT, will occur next week.

NOAA-15 rev 13264 / F at 0642Z on December 1: G3A short; data short
due to the IMUX stopping recording prematurely.  88% of the expected
data recovered.  A second attempt will be made during day shift.

POES operations were nominal over the past 24 hours.

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

The next GOES-8 North-South stationkeeping maneuver is planned for Tuesday,
December 12, with backup dates of Thursday, December 14, and Tuesday, December
19.  The maneuver is planned to start at 1753Z with a thruster firing
duration of 18 minutes and 15 seconds.

GOES-9 and GOES-11 remain in the Z-Axis Precession (ZAP) storage mode
at ~105 deg W.

GOES operations were nominal over the past 24 hours.

 

Questions or comments to [email protected]

SpaceRef staff editor.